OpenStudy (anonymous):

What is the domain of this. I got x cant equal 3 and -3.

OpenStudy (anonymous):

umm i cant do the equation, but here it is. I'm trying my best. y=x+3/4-sqr (x^2-9)

jimthompson5910 (jim_thompson5910):

The equation is \[\large y = \frac{x+3}{4-\sqrt{x^2-9}}\] right?

OpenStudy (anonymous):

yea, it doesnt show as an equation when i write it on here tho. but i think u got it right

jimthompson5910 (jim_thompson5910):

so the equation given to you is in text form like that?

OpenStudy (anonymous):

no no. the system on here is showing it in text

jimthompson5910 (jim_thompson5910):

but what I wrote is what they gave you?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok you first need to solve x^2 - 9 >= 0 for x

OpenStudy (anonymous):

i did i got -3 and 3

jimthompson5910 (jim_thompson5910):

that's if you solve x^2 - 9 = 0

jimthompson5910 (jim_thompson5910):

it's slightly different solving x^2 - 9 >= 0 , but -3 and 3 are involved

jimthompson5910 (jim_thompson5910):

a visual way to solve x^2 - 9 >= 0 is to look at the graph of x^2 - 9 and note where it is above the x axis well this is for the intervals (-infinity, -3) and (3, infinity) and you use a graph to see this therefore, the solution for x^2 - 9 >= 0 in interval notation is (-infinity, -3) U (3, infinity)

jimthompson5910 (jim_thompson5910):

the last thing you need to do is make sure the denominator is not zero (since you cannot divide by zero)

OpenStudy (anonymous):

the answer also gives me 5 and -5

jimthompson5910 (jim_thompson5910):

you need to solve \[\large 4-\sqrt{x^2-9}=0\] for x

jimthompson5910 (jim_thompson5910):

hopefully you can see how solving that leads you to x = 5 or x = -5 these two values must be excluded from the domain (to avoid division by zero)

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

yw